consumption=100+0.8Yd

investment= 150-16i

government expenditure= 100

taxes=0.25Y

DD for money=0.2Y-2i

money supply=300

price level=2

determine equilibrium level of income and rate of interest


1
Expert's answer
2021-05-13T17:36:18-0400

In goods market,

"Y = C + I + G"

"Y = 100 + 0.8(Y - 0.25Y) + 150 - 16i + 100 [\space as\space Yd = Y - T]"

"Y = 350 + 0.8 \times 0.75Y - 16i"

"Y = 350 + 0.6Y - 16i"

"0.4Y = 350 - 16i"

"Y = 875 - 40i.................. (IS\space curve)"


In money market, "Md = \frac{Ms}{P}"

"0.2Y - 2i = \frac{300}{2}"

"0.2Y - 2i = 150"

0.2Y = 150 + 2i

"Y = 750 + 10i..........(LM \space curve)"


In equilibrium, IS = LM.

"875 - 40i = 750 + 10i"

"50i = 125"

"i = 2.5"

"Y = 750 + 10 x 2.5 = 750 + 25 = 775"